OpenStudy (anonymous):

(HELP PLS?)An equilateral triangle has an apothem of 5 cm. Find the area of the triangle to the nearest square centimeter. B.130 cm C.188 cm

OpenStudy (anonymous):

Ive narrowed it down to B or C. Please help??

OpenStudy (anonymous):



OpenStudy (anonymous):

@ganeshie8 Could you help me with this? o:

OpenStudy (anonymous):

I keep getting the wrong answer for this.. i dont know how to get B or C as my answer. :S

OpenStudy (jack1):

if your distance from center mass to middle of side is 5 and its an equilateral triangle... could you use a = r cos (180/n) to find the radius then h^2 = a^2 + b^2 to find half the base length then double it to find the full base length then times that by double the apothem to find your vertical height and multiply out to find your area...?

OpenStudy (anonymous):

130?

OpenStudy (jack1):

hang on, calculating

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Was i right? or wrong?

OpenStudy (jack1):

a = r cos (180/n) a = apothem r = radius (to the corners n = number of sides so r = 10

OpenStudy (anonymous):

so i was wrong??

OpenStudy (jack1):

so side length = 17.3 verticle height = 15 so area = 129.9

OpenStudy (jack1):

u were right

OpenStudy (anonymous):

ty!!

OpenStudy (jack1):

yw